Circular No. 8180

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4641 SAGITTARII M. P. Rupen, A. J. Mioduszewski, and V. Dhawan, National Radio

Astronomy Observatory, report the detection of renewed radio activity in the x-ray transient and blackhole candidate V4641 Sgr (= SAX J1819.3– 2525). Very Large Array (VLA) 4.86-GHz observations on July 31 UT showed no source (rms noise level 0.1 mJy/beam), while data taken on Aug. 6 gave strong detections with roughly linear declines at both 4.86 (over 14 min, from 664 ± 2 mJy at Aug. 6.142 to 621 ± 2 mJy at Aug. 6.151) and 8.46 GHz (over 4 min, from 569 ± 2 mJy at Aug. 6.152 to 545 ± 2 mJy at Aug. 6.155). Extrapolating linear fits to both decays gives a spectral index, α (S_{ν} going as ν^{α}), of ∼ −0.14 at Aug. 6.1515, suggesting optically thick synchrotron radiation. Further VLA observations on Aug. 8, 9, and 10 (lasting from a few min to 1.5 hr) show a much fainter source, with flux densities ranging from 2 to 8 mJy at 4.86 GHz, and a consistent, optically thin spectral index of ∼ −0.6 between 4.86 and 8.46 GHz. These data seem to be sampling a series of radio flares, as the flux density changes abruptly from day to day and within each run; on Aug. 10, the flux density declines by a factor of two over ∼ 90 min. VLA observations on Aug. 11 give a marginal detection at 8.46 GHz of 0.34 ± 0.08 mJy and a non-detection at 4.86 GHz (3σ upper limit of 0.71 mJy). This behavior is reminiscent of that in 2002 (IAUC 7906, 7908, 7928), in which an initial very strong flare was followed by two months of continuing activity, characterized by rapid low-level flaring at optical, x-ray, and radio wavelengths. The detections, made with the VLA in its most extended (A) configuration, provide an improved radio position for the source: α = 18^{h}19^{m}21^{s}.634, δ = −25^{o}24′25′′.84 [equinox 2000.0; via the nearby astrometric calibrator ICRF J182057.8–252812 (Ma et al. 1998, A.J. 116, 516); uncertainty dominated by systematic errors (rms perhaps as large as ± 0′′.030)]. There is no sign of any extension in these images, which have a typical beam size at 8.46 GHz of 0′′.45 × 0′′.20 (FWHM of fitted Gaussian), oriented roughly north/south. D. Campbell-Wilson and R. Hunstead (University of Sydney) further report a 4.5 ± 0.8 mJy radio detection with the Molonglo Observatory Synthesis Telescope at 843 MHz, in an 8-hr observation centered on Aug. 8.465 — consistent with the VLA observations (above). Further radio observations are planned, and observations at other wavelengths are urgently requested. Simultaneous multi-wavelength observations are particularly valuable; a listing of known future observations is given at

http://www.aoc.nrao.edu/~mrupen/XRT/V4641Sgr/v4641sgr.shtml.

2003 August 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green